# Even-order optical harmonics generated from centrosymmetric-material metasurfaces

### v1

@@ Abstract
In non-conductive, bulk centrosymmetric media, even-order nonlinear optical processes can be triggered by several factors, including surface nonlinearities Timbrell et al. (2018), strong external fields leading to inversion symmetry breaking Bouhelier et al. (2003); Timurdogan et al. (2017); Lee et al. (2019), as well as by applying an external strain Schriever et al. (2010) or engineering internal strains by an additional overlayer Cazzanelli et al. (2012); Zhao et al. (2022). In the presence of free electrons in the conduction band, even-order nonlinearities may be enabled by the magnetic component of the Lorentz force Guo et al. (2019); Yang et al. (2021). Observations of second-harmonic generation (SHG) were made in various nanoscale systems made of silicon (centrosymmetric material), including membranes Scalora et al. (2019) often enhanced by Fabry-Perot resonances Hallman et al. (2023), spherical nanoparticles Makarov et al. (2017) and nanowires Wiecha et al. (2015, 2016) enhanced by Mie resonances, photonic crystal cavities Galli et al. (2010) as well as metasurfaces Bar-David and Levy (2019); Xu et al. (2021).
@@
The silicon metasurfaces are designed and fabricated to support both GMR and BIC resonances at the pump wavelength in the near-IR spectral range, with a-Si on glass and c-Si on sapphire. A schematic of the metasurfaces is shown in Fig. 1. They are composed of a square lattice of asymmetric meta-atoms consisting of bar dimers with different widths. The metasurfaces are pumped in the region from 1.4 $\mu$m to 1.75 $\mu$m, with harmonics generated in the near-infrared and visible ranges. The inset provides the scanning electron microscope (SEM) image of c-Si metasurface with the unit cell highlighted. The geometrical parameters of the unit cell are $L$=630 nm, $S$=380 nm, $w_{2}$ = 260 nm and $w_{1}$ is varied from 180 nm to 260 nm. Parameter $\Delta w=w_{2}-w_{1}$ is considered as an asymmetry parameter.
@@
Figure 2: Optical properties of c-Si metasurfaces. (a) Transmission spectrum of the metasurface with $\Delta w=60$ nm. The inset shows an SEM image of the unit cell. (b) Transmission spectra depending on the change of the asymmetric parameter $\Delta$ from an asymmetric case of 80 nm to a symmetric case of 0 nm. The black dashed line corresponds to the metasurface with $\Delta w=60$ nm.
@@
First, we characterize the linear response by measuring transmission with polarization along the long side of the bars unit cell. The transmission spectrum of the c-Si metasurface has two Fano-shaped features corresponding to GMR and BIC resonances. Figure 2a shows the transmission spectrum for the metasurfaces with an asymmetric parameter of 60 nm. The inset shows the SEM image of the metasurface unit cell. Decreasing the asymmetry parameter causes both supporting resonances to display a redshift. Figure 2b shows the transmission spectra for c-Si metasurfaces with asymmetry parameters ranging from 0 nm to 80 nm. The BIC resonance becomes less visible with decreasing $\Delta w$, whereas the GMR mode does not change in intensity. In the case of a-Si, the spectral position of the modes is shifted to shorter wavelengths, but both modes are preserved (see Figure S1 in not ). The nature of BIC and GMR modes is confirmed with their behaviour for $\Delta w=0$. The BIC mode disappears from the spectrum, while the linewidth of the GMR mode barely changes with the change of $\Delta w$.
@@
For c-Si metasurfaces the positions of the enhancement are shifted compared to the a-Si metasurfaces, similar to the linear spectra features shift. Figures 4(d,e) demonstrate the enhancement from c-Si metasurfaces depending on asymmetry parameter and pump wavelength. In contrast with a-Si metasurfaces, both modes are clearly distinguishable for SHG and THG enhancement maps. SH and TH signal enhancements are about three orders of magnitude. SH enhancement is compared to the best case scenario where the wave is incident at approximately 70 degrees  Hallman et al. (2023). A decrease in the asymmetry parameter causes the decrease of linewidth for the BIC mode and its eventual disappearance from the spectrum for the fully symmetric case. The enhancement of SHG and THG are summarized in Figure 4f by the orange and green circles, respectively. Remarkably, whereas enhancement for THG is higher for the c-Si metasurfaces and SHG is comparable due to the better quality of the structure, the absolute values of harmonics are bigger for a-Si.
@@
In summary, we have studied optical harmonic generation from metasurfaces made of amorphous and crystalline silicon. We have observed the generation of the second and third harmonics enhanced by at least three orders of magnitude via GMR and BIC resonances in the near-IR range. This enabled significant generation of the fourth harmonics at the BIC resonance. We have observed the deviation in the power laws for both even harmonics at the BIC resonance. We believe that our results reveal that dielectric metasurfaces made of centrosymmetric materials can exhibit generally unexpectedly strong even-order nonlinear effects when they are empowered by high-$Q$ optical resonances associated with the bound sytates in the continuum.
@@ Acknowledgements.
The authors thank C. Cojocaru, S. Mayer, N.C. Panoiu, J. Trull, and M.A. Vincenti for useful discussions and suggestions. This work was supported by the Australian Research Council (Grants DP210101292 and DE210100679), the International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (contract FA520923C0023), and National Natural Science Foundation of China (Grants 12334016, 12025402, and 12261131500).

### v2

@@ Abstract
In non-conductive, bulk centrosymmetric media, even-order nonlinear optical processes can be triggered by several factors, including surface nonlinearities Timbrell et al. (2018), strong external fields leading to inversion symmetry breaking Bouhelier et al. (2003); Timurdogan et al. (2017); Lee et al. (2019), as well as by applying an external strain Schriever et al. (2010) or engineering internal strains by an additional overlayer Cazzanelli et al. (2012); Zhao et al. (2022). In the presence of free electrons in the conduction band, even-order nonlinearities may be enabled by the magnetic component of the Lorentz force Guo et al. (2019); Yang et al. (2021). Observations of second-harmonic generation (SHG) were made in various nanoscale systems made of silicon (centrosymmetric material), including membranes Scalora et al. (2019) often enhanced by Fabry-Perot resonances Hallman et al. (2023), spherical nanoparticles Makarov et al. (2017) and nanowires Wiecha et al. (2015, 2016) enhanced by Mie resonances, photonic crystal cavities Galli et al. (2010) as well as metasurfaces Bar-David and Levy (2019); Xu et al. (2021); Wang et al. (2024).
@@
Figure 2: Optical properties of c-Si metasurfaces. The numerically calculated (a) and measured (c) transmission spectrum of the metasurface with $\Delta w=60$ nm. The insets show the near-field distribution of the GMR and BIC modes, and an SEM image of the unit cell. Calculated (b) and measured (d) transmission spectra depending on the change of the asymmetric parameter $\Delta$ from an asymmetric case of 80 nm to a symmetric case of 0 nm. The black dashed lines correspond to the metasurface with $\Delta w=60$ nm.
The silicon metasurfaces are designed and fabricated to support both GMR and BIC resonances at the pump wavelength in the near-IR spectral range, with a-Si on glass and c-Si on sapphire. A schematic of the metasurfaces is shown in Fig. 1. They are composed of a square lattice of asymmetric meta-atoms consisting of bar dimers with different widths. The metasurfaces are pumped in the region from 1.4 $\mu$m to 1.75 $\mu$m, with harmonics generated in the near-infrared and visible ranges. The inset provides the scanning electron microscope (SEM) image of c-Si metasurface with the unit cell highlighted. The geometrical parameters of the unit cell are bar height $h$=600 nm, $P$=830 nm, $L$=630 nm, $S$=380 nm, $w_{2}$ = 260 nm and $w_{1}$ is varied from 180 nm to 260 nm. Parameter $\Delta w=w_{2}-w_{1}$ is considered as an asymmetry parameter.
First, we characterize the linear transmission of c-Si metasurfaces in numerical calculations for incident polarization along the long side of the bars. We use the finite-element-method solver in COMSOL Multiphysics in the frequency domain. All calculations were realized for a metasurface placed on a semi-infinite substrate surrounded by a perfectly matched layer mimicking an infinite region. The simulation area is the unit cell extended to an infinite metasurface by using the Bloch boundary conditions. The material properties are taken from Ref. Li (1980). The transmission spectrum for $\Delta w=60$ nm in the wavelength range from 1400 to 1800 nm and transmission map for $\Delta w$ from 00 nm to $90$ nm are shown in Figs. 2a,b, respectively. The resonant features can be associated with three distinctive resonances, that have wavelengths of $1580$ nm, $1730$ nm and $1780$ nm for $\Delta w=60$ nm, respectively. We focus on two modes at $1580$ nm and $1730$ nm, and identify their nature via calculating near-field distributions in the eigenmode solver in COMSOL Multiphysics, shown in the inset of the Fig. 2a for $\Delta w=60$ nm. We can see that the mode at $1730$ nm is a magnetic dipolar BIC resonance because of its distinctive asymmetric field distribution, and the mode at $1580$ nm is a magnetic dipolar GMR mode. In both modes, the volume-averaged enhancement of the local field intensity of approximately 7 times is observed (see Figure S1 in not ).
@@
We then measure the transmission of the c-Si metasurfaces with polarization along the long side of the bars unit cell. The transmission spectrum has two Fano-shaped features corresponding to GMR and BIC resonances. Figure 2c shows the transmission spectrum for the metasurfaces with $\Delta w=60$ nm. The inset shows the SEM image of the metasurface unit cell. Decreasing the asymmetry parameter causes both supporting resonances to display a red shift. Figure 2d shows the transmission spectra for c-Si metasurfaces with asymmetry parameters ranging from 0 nm to 80 nm. The BIC resonance becomes less visible with decreasing $\Delta w$, whereas the GMR mode does not change in intensity. The behaviour obtained has a good agreement with theoretical prediction, which is shown in Fig. 2b. In the case of a-Si, the spectral position of the modes is shifted to shorter wavelengths, but both modes are preserved (see Figure S2 in not ).
The nature of BIC and GMR modes is numerically and experimentally confirmed with their behaviour for $\Delta w=0$. More specific, the BIC mode disappears from the spectrum, while the linewidth of the GMR mode barely changes with the change of $\Delta w$, as can be seen in Figs. 2b,d. We note that the measured resonant positions are red shifted with respect to the simulation results, and the third mode observed in numerical transmission data is not visible in the experiments. These differences can be attributed to the tolerances of fabrication and design. Also, the mode linewidth appears to be larger in the experiment. This is due to the excitation source in experiment deviating from a plane wave due to the focusing objective and a range of modes with oblique in-plane k-vectors is excited in the vicinity of the normal incidence.
@@
For c-Si metasurfaces the positions of the enhancement are shifted compared to the a-Si metasurfaces, similar to the linear spectra features shift. Figures 4(d,e) demonstrate the enhancement from c-Si metasurfaces depending on asymmetry parameter and pump wavelength. In contrast with a-Si metasurfaces, both modes are clearly distinguishable for SHG and THG enhancement maps. SH and TH signal enhancements are about three orders of magnitude. SH enhancement is compared to the best case scenario where the wave is incident at approximately 70 degrees  Hallman et al. (2023). A decrease in the asymmetry parameter causes the decrease of linewidth for the BIC mode and its eventual disappearance from the spectrum for the fully symmetric case. The enhancement of SHG and THG are summarized in Figure 4f by the orange and green circles, respectively. We note that meanwhile, the SHG enhancement is in the same order of magnitude for a-Si and c-Si metasurfaces, and the SHG signal from the c-Si metasurfaces is one order of magnitude lower than the SHG from the a-Si metasurface (see Figure S6 in not )
@@
In summary, we have studied optical harmonic generation from metasurfaces made of amorphous and crystalline silicon. We have observed the generation of the second and third harmonics enhanced by at least three orders of magnitude via GMR and BIC resonances in the near-IR range. This enabled significant generation of the fourth harmonics at the BIC resonance. We have observed the deviation in the power laws for both even harmonics at the BIC resonance. We believe that our results reveal that dielectric metasurfaces made of centrosymmetric materials can exhibit generally unexpectedly strong even-order nonlinear effects when they are empowered by high-$Q$ optical resonances associated with the bound states in the continuum.
@@ Acknowledgements.
The authors thank C. Cojocaru, S.A. Maier, N.C. Panoiu, J. Trull, and M.A. Vincenti for useful discussions and suggestions. This work was supported by the Australian Research Council (Grants DP210101292 and DE210100679), the International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (contract FA520923C0023), and National Natural Science Foundation of China (Grants 12334016, 12025402, and 12261131500).